# Equilibration and convected limit in 2D-1D corotational Oldroyd’s fluid-structure interaction

### v1

@@ 1. Introduction
is the vorticity tensor of the solvent or the antisymmetric gradient of the solvent’s velocity. To obtain a macroscopic description for the solute subcomponent from the mesoscopic description above, we consider a tensor-value average and a scalar average given by the stress tensor $\widehat{\mathbb{T}}:(t,\mathbf{x})\in I\times\Omega_{\eta}\mapsto\widehat{\mathbb{T}}(t,\mathbf{x})\in\mathbb{R}^{2\times 2}$ of the solute and the polymer number density $\rho:(t,\mathbf{x})\in I\times\Omega_{\eta}\mapsto\rho(t,\mathbf{x})\in\mathbb{R}$ which we assume to be a constant since only two monomers are at play. These averages are precisely given by
@@ 1.1. Setup
is the Hanzawa transform with inverse $\bm{\Psi}_{-\eta(t)}$ and where for a point $\mathbf{x}$ in the neighbourhood of $\partial\Omega$, the vector $\mathbf{n}(y(\mathbf{x}))$ is the unit normal at the point $y(\mathbf{x})=\mathrm{argmin}_{y\in\omega}|\mathbf{x}-\bm{\varphi}(y)|$. Also, $s(\mathbf{x})=(\mathbf{x}-\bm{\varphi}(y(\mathbf{x})))\cdot\mathbf{n}(y(\mathbf{x}))$ and $\phi\in C^{\infty}(\mathbb{R})$ is a cut-off function that is $\phi\equiv 0$ in the neighbourhood of $-L$ and $\phi\equiv 1$ in the neighbourhood of 00. Note that $\bm{\Psi}_{\eta(t)}(\mathbf{x})$ can be rewritten as
@@
$\eta(0,\cdot)=\eta_{0}(\cdot),\qquad\partial_{t}\eta(0,\cdot)=\eta_{\star}(\cdot)$ in $\text{in }\omega,$ (1.9)
$\mathbf{u}(0,\cdot)=\mathbf{u}_{0}(\cdot)$ in $\text{in }\Omega_{\eta_{0}},$ (1.10)
$\mathbb{T}(0,\cdot)=\mathbb{T}_{0}(\cdot)$ in $\text{in }\Omega_{\eta_{0}},$ (1.11)
$\mathbf{n}_{\eta}\cdot\nabla_{\mathbf{x}}\mathbb{T}=0$ on $\text{on }I\times\partial\Omega_{\eta}.$ (1.12)
@@
$\mathbf{u}\circ\bm{\varphi}_{\eta}=(\partial_{t}\eta)\mathbf{n}$ on $\text{on }I\times\omega$ (1.13)
@@
The analysis of polymeric fluid-structure interaction problems was initiated recently in [8] where the authors showed the existence of weak solutions to a dilute solute-solvent-structure mutually coupled system. This system consisted of the $3$-D noncorotational Fokker–Planck equation (1.2) (where $\nabla_{\mathbf{x}}\mathbf{u}$ replaces $\mathbb{W}(\nabla_{\mathbf{x}}\mathbf{u})$) for the mesoscopic description of the solute, the $3$-D incompressible Navier–Stokes equation (1.5) and (1.6) giving the macroscopic description of the solvent, and with a $2$-D structure modeled by a shell equation (1.7) of Koiter type (where the term $-\gamma\partial_{t}\partial_{y}^{2}\eta+\partial_{y}^{4}\eta$ is replaced by the gradient of the so-called Koiter energy). Uniqueness is unknown for this system but the solutions exist until potential degeneracies occur with the Koiter energy or with the structure deformation. When the $2$-D Koiter shell in [8] is replaced by the $2$-D viscoelastic shell equation (1.7), the extension to the existence of a unique local-in-time strong solution was then shown in [11]. Note that for fixed spatial domains subject to periodic boundary conditions, one can construct solutions that are spatially more regular [9] than strong solutions. The corresponding result for the system with a structure displacement remains an interesting open problem even in lower dimensions.
@@ 1.2. Concepts of solution and main results
The difference in the rate for $\mathbf{u}$ (i.e. the extra term $-|\Omega_{\eta}|^{-1/2}$) compared to that established in [14, Theorem 5.1] comes from the fact that the velocity field we consider does not have zero-trace. Indeed, since $\mathbf{u}$ satisfies $\mathbf{u}\circ\bm{\varphi}_{\eta}=(\partial_{t}\eta)\mathbf{n}$ on $\omega$, we obtain the extra rate for the shell velocity $\partial_{t}\eta$.
@@
$\zeta(0,\cdot)=\zeta_{0}(\cdot),\qquad\partial_{t}\zeta(0,\cdot)=\zeta_{\star}(\cdot)$ in $\text{in }\omega,$ (1.19)
$\mathbf{v}(0,\cdot)=\mathbf{v}_{0}(\cdot)$ in $\text{in }\Omega_{\zeta_{0}},$ (1.20)
$\mathbb{U}(0,\cdot)=\mathbb{U}_{0}(\cdot)$ in $\text{in }\Omega_{\zeta_{0}},$ (1.21)
@@
$\mathbf{v}\circ\bm{\varphi}_{\zeta}=(\partial_{t}\zeta)\mathbf{n}$ on $\text{on }I\times\omega$ (1.22)
at the flexible part of the boundary with normal $\mathbf{n}$. We are interested in weak solutions of (1.15)–(1.22) where weak solutions are defined as follows:
@@
A superficial disadvantage of the system (1.15)-(1.18) without any form of diffusion or damping is that the regularisation effect due to those two terms is lost. Consequently, the solute component $\mathbb{U}$ of its weak solution is less regular than the solute component $\mathbb{T}$ of the weak solution of (1.5)-(1.8), c.f. [23]. Nevertheless, because $\mathbb{U}$ possesses transport and corotational properties, any weak solution automatically regularises in the sense that it becomes essentially bounded in both space and time provided the initial conditions are equally bounded in space. Indeed, let $(\zeta,\mathbf{v},\mathbb{U})$ be a weak solution of (1.15)–(1.22) with dataset $(\mathbb{U}_{0},\mathbf{v}_{0},\zeta_{0},\zeta_{\star})$ satisfying (LABEL:mainDataForAll) and $\mathbb{U}_{0}\in L^{\infty}(\Omega_{\zeta_{0}})$. Although Equation (1.18) looks complicated, we can show that it conserves all $L^{p}$-norms uniformly in $p\in[1,\infty)$ and thus, it is essentially bounded in spacetime. To see the formal details, we test (1.18) with $\mathbb{U}^{q-1}$, $q\in[1,\infty)$ and use Proposition 2.1. We obtain
@@
Our second main result will now seek to establish a relationship between strong solutions of (1.5)-(1.13) and essentially bounded weak solutions of (1.15)–(1.22). The precise statement is given as follows.
@@
An immediate consequence of Theorem 1.8 is a weak-strong uniqueness results. Generally speaking, such a result states that a weak solution is unique in the class of a strong solution. This will usually require that both solutions solves the same equation and may also require additionally regularity of either solution or their data. See for instant [10, 13, 16, 30]. In our setting, however, (1.5)–(1.13) and (1.15)–(1.22) are two different systems, albeit related, with their own notion of a solution. Nevertheless, we present a ‘limiting’ weak-strong uniqueness result which states that as $\varepsilon\rightarrow 0$, the essentially bounded weak solution to (1.15)–(1.22) will be unique in the class of the strong solution to (1.5)–(1.13). Here, we do not require additional regularity assumption on their respective dataset. The precise statement is:
@@ 2. Preliminaries
For two non-negative quantities $F$ and $G$, we write $F\lesssim G$ if there is a constant $c>0$ such that $F\leq c\,G$. If $F\lesssim G$ and $G\lesssim F$ both hold, we use the notation $F\sim G$. The anti-symmetric gradient of a vector $\mathbf{f}\in\mathbb{R}^{d}$ is denoted by $\mathbb{W}(\nabla_{\mathbf{x}}\mathbf{f})=\tfrac{1}{2}(\nabla_{\mathbf{x}}\mathbf{f}-(\nabla_{\mathbf{x}}\mathbf{f})^{\top})$ and the scalar matrix product of the matrices $\mathbb{A}=(a_{ij})_{i,j=1}^{d}$ and $\mathbb{B}=(b_{ij})_{i,j=1}^{d}$ is denoted by $\mathbb{A}:\mathbb{B}=\sum_{ij}a_{ij}b_{ij}$. The symbol ||$\cdot$|| ⋅ | may be used in four different contexts. For a scalar function $f\in\mathbb{R}$, $|f|$ denotes the absolute value of $f$. For a vector $\mathbf{f}\in\mathbb{R}^{d}$, $|\mathbf{f}|$ denotes the Euclidean norm of $\mathbf{f}$. For a square matrix $\mathbb{F}\in\mathbb{R}^{d\times d}$, $|\mathbb{F}|$ shall denote the Frobenius norm $\sqrt{\mathrm{trace}(\mathbb{F}^{T}\mathbb{F})}$. Also, if $S\subseteq\mathbb{R}^{d}$ is a (sub)set, then $|S|$ is the $d$-dimensional Lebesgue measure of $S$. Since we only consider functions on $\omega\subset\mathbb{R}$ with periodic boundary conditions and zero mean values, we have the following equivalences
@@
$L^{p}(I;L^{r}(\Omega_{\eta}))$ :=$:=\Big{\{}v\in L^{1}(I\times\Omega_{\eta}):\begin{subarray}{c}v(t,\cdot)\in L^{r}(\Omega_{\eta(t)})\,\,\text{for a.e. }t,\\ \|v(t,\cdot)\|_{L^{r}(\Omega_{\eta(t)})}\in L^{p}(I)\end{subarray}\Big{\}},$
$L^{p}(I;W^{1,r}(\Omega_{\eta}))$ :=$:=\big{\{}v\in L^{p}(I;L^{r}(\Omega_{\eta})):\,\,\nabla_{\mathbf{x}}v\in L^{p}(I;L^{r}(\Omega_{\eta}))\big{\}}.$
@@ 4.1. Estimate for the solvent-structure interaction
[…]
@@ 4.2. Estimate for the solute
We now note that due to Ladyzhenskaya’s inequality and the fact the term $\|\nabla_{\mathbf{x}}\overline{\mathbb{T}}(t)\|_{L^{2}(\Omega_{\zeta})}$ is essentially bounded in time, we obtain
@@
×$\times\|\nabla_{\mathbf{x}}\overline{\mathbb{T}}\|_{L^{2}(\Omega_{\zeta})}^{1/2}\|\overline{\mathbb{T}}\|_{W^{2,2}(\Omega_{\zeta})}^{1/2}\big{)}\,\mathrm{d}t^{\prime}$
@@
$\quad$-- $\tfrac{1}{2}\Big{\{}(\overline{\mathbb{T}}-\mathbb{U})\nabla_{\mathbf{x}}(\overline{\mathbf{u}}-\mathbf{v}):\mathbb{U}-(\overline{\mathbb{T}}-\mathbb{U})(\nabla_{\mathbf{x}}(\overline{\mathbf{u}}-\mathbf{v}))^{\top}:\mathbb{U}\Big{\}}$

### v2

@@ 1. Introduction
is the vorticity tensor of the solvent or the antisymmetric gradient of the solvent’s velocity. To obtain a macroscopic description for the solute subcomponent from the mesoscopic description above, we consider a tensor-valued average and a scalar average given by the stress tensor $\widehat{\mathbb{T}}:(t,\mathbf{x})\in I\times\Omega_{\eta}\mapsto\widehat{\mathbb{T}}(t,\mathbf{x})\in\mathbb{R}^{2\times 2}$ of the solute and the polymer number density $\rho:(t,\mathbf{x})\in I\times\Omega_{\eta}\mapsto\rho(t,\mathbf{x})\in\mathbb{R}$ which we assume to be a constant since only two monomers are at play. These averages are precisely given by
@@ 1.1. Setup
is the Hanzawa transform with inverse $\bm{\Psi}_{-\eta(t)}$ and where for a point $\mathbf{x}$ in the neighbourhood of $\partial\Omega$, the vector $\mathbf{n}(y(\mathbf{x}))$ is the unit normal at the point $y(\mathbf{x})=\mathrm{argmin}_{y\in\omega}|\mathbf{x}-\bm{\varphi}(y)|$. Also, $s(\mathbf{x})=(\mathbf{x}-\bm{\varphi}(y(\mathbf{x})))\cdot\mathbf{n}(y(\mathbf{x}))$ and $\phi\in C^{\infty}(\mathbb{R})$ is a cut-off function that is $\phi\equiv 0$ in the neighbourhood of L-L and $\phi\equiv 1$ in the neighbourhood of 0. Note that $\bm{\Psi}_{\eta(t)}(\mathbf{x})$ can be rewritten as
@@
$\eta(0,\cdot)=\eta_{0}(\cdot),\qquad\partial_{t}\eta(0,\cdot)=\eta_{\star}(\cdot)$ $\text{in }\omega,$ (1.9)
$\mathbf{u}(0,\cdot)=\mathbf{u}_{0}(\cdot)$ $\text{in }\Omega_{\eta_{0}},$ (1.10)
$\mathbb{T}(0,\cdot)=\mathbb{T}_{0}(\cdot)$ $\text{in }\Omega_{\eta_{0}},$ (1.11)
$\mathbf{n}_{\eta}\cdot\nabla_{\mathbf{x}}\mathbb{T}=0$ $\text{on }I\times\partial\Omega_{\eta}.$ (1.12)
@@
$\mathbf{u}\circ\bm{\varphi}_{\eta}=(\partial_{t}\eta)\mathbf{n}$ $\text{on }I\times\omega$ (1.13)
@@
The analysis of polymeric fluid-structure interaction problems was initiated recently in [8] where the authors showed the existence of weak solutions to a dilute solute-solvent-structure mutually coupled system. This system consisted of the $3$-D noncorotational Fokker–Planck equation (1.2) (where $\nabla_{\mathbf{x}}\mathbf{u}$ replaces $\mathbb{W}(\nabla_{\mathbf{x}}\mathbf{u})$) for the mesoscopic description of the solute, the $3$-D incompressible Navier–Stokes equation (1.5) and (1.6) giving the macroscopic description of the solvent, and with a $2$-D structure modeled by a shell equation (1.7) of Koiter type (where the term $\gamma\partial_{t}\partial_{y}^{2}\eta+\partial_{y}^{4}\eta$ is replaced by the gradient of the so-called Koiter energy). Uniqueness is unknown for this system but the solutions exist until potential degeneracies occur with the Koiter energy or with the structure deformation. When the $2$-D Koiter shell in [8] is replaced by the $2$-D viscoelastic shell equation (1.7), the extension to the existence of a unique local-in-time strong solution was then shown in [11]. Note that for fixed spatial domains subject to periodic boundary conditions, one can construct solutions that are spatially more regular [9] than strong solutions. The corresponding result for the system with a structure displacement remains an interesting open problem even in lower dimensions.
@@ 1.2. Concepts of solution and main results
The difference in the rate for $\mathbf{u}$ (i.e. the extra term $|\Omega_{\eta}|^{-1/2}$) compared to that established in [14, Theorem 5.1] comes from the fact that the velocity field we consider does not have zero-trace. Indeed, since $\mathbf{u}$ satisfies $\mathbf{u}\circ\bm{\varphi}_{\eta}=(\partial_{t}\eta)\mathbf{n}$ on $\omega$, we obtain the extra rate for the shell velocity $\partial_{t}\eta$.
@@
$\zeta(0,\cdot)=\zeta_{0}(\cdot),\qquad\partial_{t}\zeta(0,\cdot)=\zeta_{\star}(\cdot)$ $\text{in }\omega,$ (1.19)
$\mathbf{v}(0,\cdot)=\mathbf{v}_{0}(\cdot)$ $\text{in }\Omega_{\zeta_{0}},$ (1.20)
$\mathbb{U}(0,\cdot)=\mathbb{U}_{0}(\cdot)$ $\text{in }\Omega_{\zeta_{0}},$ (1.21)
@@
$\mathbf{v}\circ\bm{\varphi}_{\zeta}=(\partial_{t}\zeta)\mathbf{n}$ $\text{on }I\times\omega$ (1.22)
at the flexible part of the boundary with normal $\mathbf{n}$. A similar system was analysed in [25, (1.13)-(1.17)] with an extra equation for the evolution of the polymer number density $\sigma$ and with an additional damping term $2(\mathbb{U}-\sigma\mathbb{I})$ in the equation for $\mathbb{U}$. For dilute polymers, monomer-to-monomer interactions are negligible and thus it is reasonable to assume $\sigma$ is a constant. In this case, there is no evolutionary equation for $\sigma$ and $\mathbb{U}=\widehat{\mathbb{U}}-\sigma\mathbb{I}$ satisfy (1.18), where
$\widehat{\mathbb{U}}(t,\mathbf{x})=\int_{B}f(t,\mathbf{x},\mathbf{q})\mathbf{q}\otimes\mathbf{q}\,\mathrm{d}\mathbf{q},\qquad\varepsilon=\frac{1}{2\lambda}=0.$
Moreover, from the purely analytical point of view, the equation for $\sigma$ which is a homogeneous transport equation does not impose any analytical difficulty that is not already captured by the equation for $\mathbb{U}$. As a result, many analysts, see for instant [12, 18, 32], elect to study the system without $\sigma$. Putting this minor difference aside, the main difference between (1.15)-(1.18) and the limit system analysed in [25, (1.13)-(1.17)] is that the former does not have the damping term $2(\mathbb{U}-\sigma\mathbb{I})$ (or just $2\mathbb{U}$ when the polymer number density is a constant) present in the latter. This term is a hyperbolic dissipation term that regularises the equation for the solute (without needing a more regular initial condition) from the mathematical point of view. From the physical point of view, this captures the system with no localised damping in the solute.
Furthermore, the purely hyperbolic equation (1.18) has ‘truly’ infinite memory since its stress never relaxes. A consequence of this is that there is no exponential decay of stress and thus, Theorem 1.2 or any variant of it is completely out of reach. Nevertheless, as pointed out earlier, our second main result will study the singular limit of the system (1.5)-(1.8) with finite memory to the system (1.15)-(1.18) with this truly infinite memory. It is worth noting that although the limit system in [25, (1.13)-(1.17)] also has infinite memory, this is weaker in the sense that it’s memory is exponentially fading but nonzero for all past times. Finally, we also note that, whereas [25, (1.13)-(1.17)] is analytically delicate yet manageable to handle, the former is extremely hard and becomes comparable in difficulty to the 3-D Euler equations when the full velocity gradient is considered in the solute equation (rather than the anti-symmetric gradient).
We are interested in weak solutions of (1.15)–(1.22) where weak solutions are defined as follows:
@@
A superficial disadvantage of the system (1.15)-(1.18) without any form of diffusion or damping is that the regularisation effect due to those two terms is lost. Consequently, the solute component $\mathbb{U}$ of its weak solution is less regular than the solute component $\mathbb{T}$ of the weak solution of (1.5)-(1.8), c.f. [23]. Nevertheless, because $\mathbb{U}$ possesses transport and corotational properties, any weak solution automatically regularises in the sense that it becomes essentially bounded in both space and time provided the initial condition is essentially bounded in space. Indeed, let $(\zeta,\mathbf{v},\mathbb{U})$ be a weak solution of (1.15)–(1.22) with dataset $(\mathbb{U}_{0},\mathbf{v}_{0},\zeta_{0},\zeta_{\star})$ satisfying (LABEL:mainDataForAll) and $\mathbb{U}_{0}\in L^{\infty}(\Omega_{\zeta_{0}})$. Although Equation (1.18) looks complicated, we can show that it conserves all $L^{p}$-norms uniformly in $p\in[1,\infty)$ and thus, it is essentially bounded in spacetime. To see the formal details, we test (1.18) with $\mathbb{U}^{q-1}$, $q\in[1,\infty)$ and use Proposition 2.1. We obtain
@@
Our second main result will now seek to establish a relationship between strong solutions of (1.5)-(1.13) whose solution has finite memory and essentially bounded weak solutions of (1.15)–(1.22) with truly infinite memory. The precise statement is given as follows.
@@
An immediate consequence of Theorem 1.8 is a weak-strong uniqueness result. Generally speaking, such a result states that a weak solution is unique in the class of a strong solution. This will usually require that both solutions solves the same equation and may also require additionally regularity of either solution or their data. See for instant [10, 13, 16, 30]. In our setting, however, (1.5)–(1.13) and (1.15)–(1.22) are two different systems, albeit related, with their own notion of a solution. Nevertheless, we present a ‘limiting’ weak-strong uniqueness result which states that as $\varepsilon\rightarrow 0$, the essentially bounded weak solution to (1.15)–(1.22) will be unique in the class of the strong solution to (1.5)–(1.13). Here, we do not require additional regularity assumption on their respective dataset. The precise statement is:
@@ 2. Preliminaries
For two non-negative quantities $F$ and $G$, we write $F\lesssim G$ if there is a constant $c>0$ such that $F\leq c\,G$. If $F\lesssim G$ and $G\lesssim F$ both hold, we use the notation $F\sim G$. The anti-symmetric gradient of a vector $\mathbf{f}\in\mathbb{R}^{d}$ is denoted by $\mathbb{W}(\nabla_{\mathbf{x}}\mathbf{f})=\tfrac{1}{2}(\nabla_{\mathbf{x}}\mathbf{f}-(\nabla_{\mathbf{x}}\mathbf{f})^{\top})$ and the scalar matrix product of the matrices $\mathbb{A}=(a_{ij})_{i,j=1}^{d}$ and $\mathbb{B}=(b_{ij})_{i,j=1}^{d}$ is denoted by $\mathbb{A}:\mathbb{B}=\sum_{ij}a_{ij}b_{ij}$. The symbol $|\cdot|$ may be used in four different contexts. For a scalar function $f\in\mathbb{R}$, $|f|$ denotes the absolute value of $f$. For a vector $\mathbf{f}\in\mathbb{R}^{d}$, $|\mathbf{f}|$ denotes the Euclidean norm of $\mathbf{f}$. For a square matrix $\mathbb{F}\in\mathbb{R}^{d\times d}$, $|\mathbb{F}|$ shall denote the Frobenius norm $\sqrt{\mathrm{trace}(\mathbb{F}^{T}\mathbb{F})}$. Also, if $S\subseteq\mathbb{R}^{d}$ is a (sub)set, then $|S|$ is the $d$-dimensional Lebesgue measure of $S$. Since we only consider functions on $\omega\subset\mathbb{R}$ with periodic boundary conditions and zero mean values, we have the following equivalences
@@
$L^{p}(I;L^{r}(\Omega_{\eta}))$ $:=\Big\{v\in L^{1}(I\times\Omega_{\eta}):\begin{subarray}{c}v(t,\cdot)\in L^{r}(\Omega_{\eta(t)})\,\,\text{for a.e. }t,\\ \|v(t,\cdot)\|_{L^{r}(\Omega_{\eta(t)})}\in L^{p}(I)\end{subarray}\Big\},$
$L^{p}(I;W^{1,r}(\Omega_{\eta}))$ $:=\big\{v\in L^{p}(I;L^{r}(\Omega_{\eta})):\,\,\nabla_{\mathbf{x}}v\in L^{p}(I;L^{r}(\Omega_{\eta}))\big\}.$
@@ 4.1. Estimate for the solvent-structure interaction
$\big($ $\|(\overline{\mathbf{u}}-\mathbf{v})(t)\|_{L^{2}(\Omega_{\zeta})}^{2}+\|\partial_{t}(\eta-\zeta)(t)\|_{L^{2}(\omega)}^{2}+\|\partial_{y}^{2}(\eta-\zeta)(t)\|_{L^{2}(\omega)}^{2}\big)$ (4.7)
@@ 4.2. Estimate for the solute
The main differences between (4.9) and [25, (4.20)] is the absence of the damping difference present on the left and side of [25, (4.20)] as well as the $\varepsilon$-parameter term $K_{6}$ in (4.9) which does not also appear in [25, (4.20)]. Since we have optimised the analysis in [25, (4.20)] to avoid utilising the damping difference, we can essentially repeat the estimates that are common in our both settings for completeness. Furthermore, the $\varepsilon$-parameter term $K_{6}$ will be fairly easy to deal with as we shall soon see. Moving on, we now note that due to Ladyzhenskaya’s inequality and the fact that the term $\|\nabla_{\mathbf{x}}\overline{\mathbb{T}}(t)\|_{L^{2}(\Omega_{\zeta})}$ is essentially bounded in time, we obtain
@@
$\times\|\nabla_{\mathbf{x}}\overline{\mathbb{T}}\|_{L^{2}(\Omega_{\zeta})}^{1/2}\|\overline{\mathbb{T}}\|_{W^{2,2}(\Omega_{\zeta})}^{1/2}\big)\,\mathrm{d}t^{\prime}$
@@
$\quad$- $\tfrac{1}{2}\Big\{(\overline{\mathbb{T}}-\mathbb{U})\nabla_{\mathbf{x}}(\overline{\mathbf{u}}-\mathbf{v}):\mathbb{U}-(\overline{\mathbb{T}}-\mathbb{U})(\nabla_{\mathbf{x}}(\overline{\mathbf{u}}-\mathbf{v}))^{\top}:\mathbb{U}\Big\}$